In an experiment of titration, 25.0 mL of potassium hydroxide solution is poured

into a conical flask. A few drops of phenolphthalein are added into the solution.

The solution in the conical flask is titrated with 1.0 M sulphuric acid.

a)What is the function of phenolphthalein in this experiment

b)Write a balanced chemical equation for the reaction.

c) In this experiment, 20.00 mL sulphuric acid needed to react with 25.0 

mL of potassium hydroxide.

i. Calculate the mole for sulphuric acid.

ii. Calculate the molarity of potassium hydroxide solution.


1
Expert's answer
2021-10-27T06:41:17-0400

a) Phenolphthalein is the indicator helping to identify the equivalence point of the titration.


b) 2KOH + H2SO4 = K2SO4 + 2H2O


c)

i) According to this equation, n (KOH) = 2 x n (H2SO4)

CM = n / V

n = CM x V

n (H2SO4) = 1.0 x 0.020 = 0.020 mol

ii) n (KOH) = 2 x n (H2SO4) = 2 x 0.020 = 0.04 mol

CM (KOH) = 0.04 / 0.025 = 1.6 M